OpenStudy (princess_dymie):

What is the value of the expression? (1+5)*2-5*2

OpenStudy (danieldjpon3):

what do u mean?

OpenStudy (danieldjpon3):

like an answer?

OpenStudy (danieldjpon3):

the answer is 2

OpenStudy (3mar):

\[2*(1+5)-5^2=2*6-5^2=12-25=-13\]

OpenStudy (danieldjpon3):

use PEMDAS

OpenStudy (3mar):

Why 2 , @danieldjpon3?

OpenStudy (danieldjpon3):

so (1+5) is 6 then times 2, which is 12. Then do 2x5 u get 10 then 12 minus 10 = 2

OpenStudy (danieldjpon3):

idk lol

OpenStudy (mathgeek289):

(1+5) * 2 - 5 * 2 = X Pemdas. 1+5 = 6. 6 * 2 - 5 * 2 = X 5 * 2 = 10 6 * 2 = 12. 12 - 10 = X 12 - 10 = 2 2 = X So, Your Answer Is 2.

OpenStudy (danieldjpon3):

See

OpenStudy (danieldjpon3):

told ya

OpenStudy (danieldjpon3):

:D

OpenStudy (3mar):

Oh sorry \[(2*6)-(5*2)=12-10=2\] Thanks, danieldjpon3.